A plane travels at x mph. There is a tailwind of 50 mph. How many hour will it take for the plane to travel 350mph?
Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)


Hi,
Assuming 350mi is the distance traveled.
D = r*t Or t = D/r
plane traveling 350mi, at x mph with a tailwind of 50 mph
t = 350mi/(x + 50mph)
Might repost and make the question more complete, so as we can best assist You.

Answer by Alan3354(69443)

350 mph is a speed, not a distance.